The bright star Betelgeuse in the constellation Orion emits the greatest intensity of radiation at a wavelength of 808 nm. Calculate the surface temperature of the star.
Answer #1

Given :

Wavelength corresponding to greatest intensity of

radiation \lambda p = 808 nm = 808* 10^-9 m

Surface temperature T =?

Surface temperature can be calculated using 'Wein's displacement law' :

\lambdap* T = b

( where b = 2.898 * 10-3​​​​​​​​​​ m- K , called as Wein's displacement constant and temperature in Kelvin)

\lambdap T = 2.898* 10-3

808* 10-9 * T = 2.898 * ​​​​​​10-3

T = 3586.63 K

Hence surface temperature of the star = T = 3586.63 K